A 800 kg car pushes a 1900 kg truck that has a dead battery. When the driver steps on the accelerator, the drive wheels of the car push against the ground with a force of 4000 N . Rolling friction can be neglected.

What is the magnitude of the force of the car on the truck?

What is the magnitude of the force of the truck on the car?
Answer #1

A.

Fct = mt*a

a = F/(m1+m2)

Fct = 1900*4000/(1900+800)

Fct = 1900*4000/2700 = 2814.81 N

B. according to newtons third law the 2 forces in this question are equal and opposite

Ftc = -2814.81 N

|Ftc| = 2814.81 N